OpenStudy (anonymous):

Eight hundred dollars is borrowed at an interest rate of 1.5% per month. Find the interest paid over a period of 5 months.

OpenStudy (anonymous):

If interest is compounded Let A(t) = account balance at end of t months. Then A(0) = 800 A(5) = 800*1.015^5 Interest = A(5) - A(0) = 800*1.015^5 - 800 = 61.83 If interest is simple Interest = 800*.015*5 = 60

OpenStudy (anonymous):

the interest is 12 hundred dollars. 12 times 5= 60